if you know the perimenter and length of the rectangle then how do you find the area and the width ?
Answer by checkley71(8403)

PERIMETER=2*LENGTH+2*WIDTH SO IF YOU KNOW THE PERIMETER AND THE LENGTH THUS
PERIMETER IS SAY 48 & THE LENGTH IS 10 THEN 48=2*10+2W OR 48=20+2W OR 28=2W OR
W=28/2 OR W=14 WHICH IS THE WIDTH
HAVING FOUND THE WIDTH & HAVING THE LENGTH WE NOW CAN FIND THE AREA
AREA=L*W OR IN THIS CASE THE A=10*14 OR A=140